# Locally Private Sampling with Public Data

### v1

@@ 5.2 Synthetic Dataset
We use synthetic data to simulate the local distribution of a single user and their public prior, $q$. Our goal is to examine how the gap between the user’s original distribution and the sampling distribution is affected by the user’s data.
In each round, the elements of $q$ are randomly sampled from the uniform distribution, $U[0,1]$, and are normalized to sum to 1. The local distribution $p$ is constructed by fixing the first element, $p^{1}$, and varying it from $1/n$ to 1, while keeping the remaining elements equal to ensure their total sum is also 1. As $p^{1}$ increases, $p$ becomes more concentrated, approximating a Dirac distribution. This allows us to observe scenarios where our approach, designed to minimize worst-case divergence, is expected to outperform the relative mollifier sampling framework due to Lemma 1.
@@ 5.3 Avezu Click Rate Prediction Dataset
Task: The goal is to predict the next website where a user is most likely to click on an ad, while protecting their private click history. Each user’s empirical distribution of their click history is treated as their individual distribution, $p$. To simulate a public prior, users are grouped based on anonymized attributes and subcategories. The average empirical distribution of click histories within each subcategory serves as the public prior, $q$, for each group.
We select the top 100 websites with the most clicks in each subcategory to predict the website most likely to receive a click from users. We keep users with at least 20 clicks on these websites and limit selection to subcategories with at least $1000$ users. This ensures each subcategory represents a sufficiently large, identifiable demographic, where public data, such as aggregate statistics, would likely be available.
Reporting Metric: We measure the worst-case ${\mathsf{TV}}$-distance between the users’ local distributions and their corresponding sampling distributions obtained using our minimax optimal approach and the relative mollifier sampling framework.
@@
Figure 3: Comparison of our private sampling method with the relative mollifier sampling framework for predicting the next website for an AdTech company to display ads on. The maximum ${\mathsf{TV}}$-distance between users’ local distributions and their corresponding sampling distributions is shown for attribute C18 and all its subcategories with $\varepsilon=12$.
5.4 MovieLens 1M Dataset
Dataset: MovieLens 1M dataset consists of 1 million ratings provided by 6,000 users who have rated at least 20 movies. The dataset includes demographic information, and we focus on age range in our analysis.
@@ 5.4 MovieLens 1M Dataset
Task: The goal is to simulate a streaming company predicting the genre of the next movie users are likely to watch based on their private past movie rating history. Each user’s local distribution, $p$, is obtained by grouping the rated movies by their primary genre, summing the ratings given to each genre, and normalizing these sums to form a probability distribution. To estimate the public prior, $q$, users are grouped by age range, and the same process is repeated within each group.
@@
Reporting Metric: We report the maximum ${\mathsf{TV}}$-distance between users’ local distributions and the corresponding sampling distributions. This comparison is made across all age groups using a privacy parameter of $\varepsilon=4$.
Result: As shown in Figure 4, our private sampling approach significantly reduces the maximum ${\mathsf{TV}}$-distance compared to the relative mollifier sampling framework. For example, in the 1818181824242424 age range, the maximum ${\mathsf{TV}}$-distance is halved from $0.62$ to less than $0.3$, demonstrating a substantial improvement in preserving the worst-off user’s distribution.
@@
Figure 4: Comparison of our private sampling method with the relative mollifier sampling framework for predicting the genre of the next movie users are likely to watch. This figure reports the maximum ${\mathsf{TV}}$-distance between users’ local distributions and their corresponding sampling distributions across age groups, with $\varepsilon=4$.
@@
In summary, our experiments consistently demonstrate that our minimax optimal private sampler significantly outperforms the state-of-the-art approach, all while maintaining the same level of privacy guarantees.
6 Conclusion
In this paper, we introduced the concept of minimax optimal private samplers paving the way for private synthetic data generation. For discrete distributions, we fully characterized the optimal utility for these samplers and provided a recursive algorithm to compute the optimal mechanism. Additionally, we showed that the optimal mechanism is universal among all $f$-divergences. The experiments validate our theoretical findings, showcasing the minimax optimality of our approach compared to the state-of-the-art locally private sampler across various scenarios.
@@ A.8 Linear Optimization for [23]
Where $\hat{p}$ is the projected distribution. This linear program is solved using the Python interface of the Gurobi optimization package using their free license. The code for this optimization is available in the code file attached to the submission.
A.9 Additional Experiments
@@
A.10 Computing Infrastructure

### v2

@@ 5.2 Synthetic Dataset
We use synthetic data to simulate the local distribution of a single user and their public prior. Our goal is to examine how the gap between the user’s original distribution and the sampling distribution is affected by the user’s data.
In each round, the elements of the public distribution, $q$, are randomly sampled from the uniform distribution, $U[0,1]$, and are normalized to sum to 1. The local distribution $p=[p^{1},\dots,p^{n}]$ is constructed by fixing the first element, $p^{1}$, and varying it from $1/n$ to 1, while keeping the remaining elements equal to ensure their total sum is also 1. As $p^{1}$ increases, $p$ becomes more concentrated, approximating a Dirac distribution. This allows us to observe scenarios where our approach, designed to minimize worst-case divergence, is expected to outperform the relative mollifier sampling framework due to Lemma 1.
@@ 5.3 Avezu Click Rate Prediction Dataset
Task: The goal is to simulate an AdTech company selecting the optimal website to display an ad where a user is most likely to click, while ensuring the protection of their private click history. Each user’s empirical distribution of their click history is treated as their individual distribution, $p$. To approximate access to a public prior, we group users based on anonymized attributes and subcategories. The average empirical distribution of click histories within each subcategory serves as the public prior, $q$, for each group.
We select the top 100 websites with the most clicks in each subcategory to predict the website most likely to receive a click from users. We keep users with at least 20 clicks on these websites and limit selection to subcategories with at least $100$ users. This ensures each subcategory represents a sufficiently large, identifiable demographic, where public data, such as aggregate statistics, would likely be available.
Reporting Metric: We calculate the ${\mathsf{TV}}$-distance between users’ local distributions and their corresponding sampling distributions, obtained using our minimax optimal approach and the relative mollifier sampling framework, and report the maximum observed ${\mathsf{TV}}$-distance across all users for each attribute and subcategory.
@@
Figure 3: Comparison of our private sampling method with the relative mollifier sampling framework for inferring the next website for an AdTech company to display ads on. For each user, we first compute the ${\mathsf{TV}}$-distance between their local distribution and the corresponding sampling distribution. The figure then reports the maximum ${\mathsf{TV}}$-distance observed within each Subcategory for $\varepsilon=12$.
5.4 MovieLens Datasets
Datasets: The MovieLens 100K and MovieLens 1M datasets contain 100,000 and 1 million ratings, respectively, provided by 1,000 and 6,000 users. Both datasets include demographic details, with our analysis focusing on age range.
Task: The goal is to simulate a streaming company’s decision-making process for recommending the next movie genre to a user based on their age demographic while protecting their privacy.
@@ 5.4 MovieLens Datasets
For each user, we construct a local distribution, $p$, by categorizing their rated movies based on the primary genre, summing the ratings within each genre, and normalizing these sums to obtain a probability distribution. With approximately 20 genres in the dataset, this results in a distribution over genre categories. To estimate the public prior, $q$, we group users from the MovieLens 100K dataset based on their age range and apply the same procedure within each group.
@@
Reporting Metric: We calculate the ${\mathsf{TV}}$-distance between the local distributions of users and their corresponding sampling distributions and report the maximum among all users. This comparison is performed in age groups with a privacy parameter of $\varepsilon=5$.
@@
Figure 4: Comparison of our private sampling method with the relative mollifier sampling framework for inferring the genre of the next movie users are likely to watch. For each user, we first compute the ${\mathsf{TV}}$-distance between their local distribution and the corresponding sampling distribution. The figure then reports the maximum ${\mathsf{TV}}$-distance observed within each age group for $\varepsilon=5$.
Result: As shown in Figure 4, our method achieves a 41% overall improvement in preserving the private distribution compared to the RM framework, demonstrating a substantial enhancement in maintaining the worst-off user’s distribution.
Additionally, we present the results for $\varepsilon=1$ to $8$, as shown in Section A.10. When considering all privacy parameters, our approach outperforms the relative mollifier sampling framework in 54.1% of cases, achieving an average improvement of 0.11 in the maximum ${\mathsf{TV}}$-distance. In contrast, the relative mollifier sampling framework outperforms our approach in 41.6% of cases, with a much smaller average improvement of 0.03 in the maximum ${\mathsf{TV}}$-distance.
@@
In summary, our experiments show that our minimax-optimal private sampler consistently outperforms the state-of-the-art approach across various scenarios, all while preserving the same level of privacy guarantees.
6 Disclosure of Funding
The work of B. Zamanlooy was supported in part by the NSERC Gradute Scholarsip. The work of M. Diaz was supported in part by the Programa de Apoyo a Proyectos de Investigación e Innovación Tecnológica (PAPIIT) under grant IN103224. The work of S. Asoodeh was supported in part by the NSERC of Canada.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes/No/Not Applicable] Yes. We have described Algorithm 1 in 4 steps in section 2 and provided the pseudo code.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes/No/Not Applicable] Yes. The time complexity analysis of Algorithm 1 is given at the end of Section 3.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes/No/Not Applicable] Yes - We have included the code with specification of all dependencies for all experiments in the supplementary material.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes/No/Not Applicable] Yes. All assumptions are stated for each theoretical claim in the main body of the paper.
    2. (b)
      Complete proofs of all theoretical results. [Yes/No/Not Applicable] Yes. These are provided in the appendix. Additionally, we provide a sketch of the proofs in the main body of the paper.
    3. (c)
      Clear explanations of any assumptions. [Yes/No/Not Applicable] Yes. All assumptions are stated for each theoretical claim in the main body of text.
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes/No/Not Applicable] Yes. https://github.com/bzamanlooy/LDP-Sampling-Public-data
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes/No/Not Applicable] No Applicable.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes/No/Not Applicable] Yes. This is provided in the main body of the text in the experiments section.
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes/No/Not Applicable] Yes. We provide it in Section A.11 and mention it in the main text.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. Yes. We have cited the creators of the datasets we used in our experiments.
    2. (b)
      The license information of the assets, if applicable. Not Applicable
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. Not Applicable
    4. (d)
      Information about consent from data providers/curators. Not Applicable
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. Not Applicable
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. Not Applicable
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. Not Applicable
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. Not Applicable
@@ A.7 Proof of Corollary 2
Where $\hat{p}$ is the projected distribution. This linear program is solved using the Python interface of the Gurobi optimization package using their free license.
A.9 Additional Experiments - Avezu Click Rate Dataset
@@
A.10 Additional Experiments - MovieLens Dataset
Figure 12: Comparison of our private sampling method with the relative mollifier sampling framework for predicting the genre of the next movie users are likely to watch. The figure reports the maximum ${\mathsf{TV}}$-distance between users’ local distributions and their corresponding sampling distributions across different age groups.
A.11 Computing Infrastructure